OpenStudy (anonymous):

A video game club charges a fixed annual membership fee of $20 and $2 per video game rented. Let f(n) represent the total annual cost of renting n video games. Which of the following functions best represents the relationship between f(n) and n if the membership was increased by $15 the next year?

OpenStudy (anonymous):

f(n) = 2n + 35 f(n) = 20n + 2 f(n) = 10n + 35 f(n) = 2n + 10

OpenStudy (anonymous):

f(n) = 2n + 35